# Quantum and Classical Communication Complexity of Permutation-Invariant Functions

### v1

@@ 3.2 Randomized and Quantum Communication Complexity Upper Bound
      $p=\frac{|x\land y|}{n-(a+b)+2|x\land y|}=\frac{1}{\frac{n-(a+b)}{|x\land y|}+2}\enspace.$
      Notice that $p$ is an increasing function with respect to $|x\land y|$. As a result, if $|x\land y|\leq c-g$, then $p\leq\frac{c-g}{m-2g}$; if $|x\land y|\geq c+g$, then $p\geq\frac{c+g}{m+2g}$. Let
      $\beta\coloneqq\frac{1}{2}\left(\frac{c+g}{m+2g}+\frac{c-g}{m-2g}\right)=\frac{cm-2g^{2}}{m^{2}-4g^{2}}\text{ and }\epsilon\coloneqq\frac{1}{2}\left(\frac{c+g}{m+2g}-\frac{c-g}{m-2g}\right)=\frac{gm}{m^{2}-4g^{2}}\enspace.$
      Since $c-g\geq 0$, we have $\beta\leq\frac{c+g}{m+2g}\leq\frac{2c}{m}$ and
      $\epsilon$ $=\frac{1}{2}\left(\frac{c+g}{m+2g}-\frac{c}{m}\right)+\frac{1}{2}\left(\frac{c}{m}-\frac{c-g}{m-2g}\right)$
      $=\frac{1}{2}\left(\frac{g(m-2c)}{m(m+2g)}+\frac{g(m-2c)}{m(m-2g)}\right)$
      $=O\left(\frac{g}{m}\right)\enspace.$
      For any $x\in\{0,1\}^{n}$, we let $S_{x}\coloneqq\left\{i:x_{i}=1\right\}$. By 3.12, Alice and Bob can sample $i$ from $S_{\overline{x}\oplus y}$ uniformly using $O(\log n)$ bits communication. Since $i\in S_{\overline{x}\oplus y}$, if $x_{i}=y_{i}=1$, then $i\in S_{x\land y}$; if $x_{i}=y_{i}=0$, then $i\notin S_{x\land y}$. By 3.11, using $O\left(\frac{\beta}{\epsilon^{2}}\right)=O\left(\frac{mc}{g^{2}}\right)=O\left(\frac{n_{1}n_{2}}{g^{2}}\right)$ samples, Alice and Bob can decide whether $p\leq\beta-\epsilon$ or $p\geq\beta+\epsilon$ with bounded error. Equivalently, Alice and Bob can distinguish $|x\land y|\leq c-g$ from $|x\land y|\geq c+g$ with bounded error. By error reduction, using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log\log n\right)$ samples, they can decide whether $|x\land y|\geq c-g$ or $|x\land y|\leq c+g$ with success probability at least $1-1/(6\log n)$. Thus, the communication complexity is $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$.
@@
      Case 3.2: $a+b=n$. Alice and Bob generate new inputs $x^{\prime}=x0$ and $y^{\prime}=y0$ (pad a zero after the original input). We know
      $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)=\mathsf{SetInc}_{a,b,c,g}^{n+1}(x^{\prime},y^{\prime})\enspace.$
@@ 4 Log-Rank Conjecture for Permutation-Invariant Functions
Theorem 1.6 states the Log-rank Conjecture for permutation-invariants functions. We argue for the lower bound (Lemma 4.1) and the upper bound (Lemma 4.2) separately.
Lemma 4.1.
Fix $n\in\mathbb{N}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be a non-trivial total permutation-invariant function. For every $a,b\in[n]$ such that $f_{a,b}$ is not a constant function, we have
@@
where $f_{a,b}$ is defined as Equation 1.
Lemma 4.2.
Fix $n\in\mathbb{N}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be a non-trivial total permutation-invariant function.
$D(f)=O\left(\max_{a,b\in[n]:f_{a,b}\text{ is not constant}}\min\left\{a,b,n-a,n-b\right\}\cdot\log n\right)\enspace,$
where $f_{a,b}$ is defined as Equation 1.
We prove Lemma 4.1 below, and postpone the proof of Lemma 4.2 to Appendix C.
Proof of Lemma 4.1.
@@
Fact 4.3 ([16], merging Corollary 6 with Lemma 4).
Fix $n\in\mathbb{N}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be defined as $f(x,y)\coloneqq D(|x\land y|)$ for some predicate $D:\left\{0,1,...,n\right\}\rightarrow\left\{-1,1\right\}$. If $t$ is the smallest integer such that $D(t)\neq D(t-1)$, then $\log\operatorname{\mathrm{rank}}(f)=\Omega\left(\log\left(\sum_{i=t}^{n}\binom{n}{i}\right)\right)$.
Fact 4.4.
Fix $n\in\mathbb{N}$. Let $\mathcal{X},\mathcal{Y}\coloneqq\left\{x\in\{0,1\}^{n}:|x|=k\right\}$, where $k\leq n/2$. Let $\operatorname{\mathsf{DISJ}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{-1,1\right\}$ and $\operatorname{\mathsf{EQ}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{-1,1\right\}$ be defined as
@@
We refer interesting readers to Section A.2 for the proof of 4.4. Now we prove the lemma by casing on the values of $a$ and $b$.
@@
$\operatorname{\mathsf{DISJ}}_{n^{\prime}}^{1}(x,y)=f_{a,b}(|x\land y|+c)=f(x^{\prime},y^{\prime})\enspace,$
@@
$x^{\prime}\coloneqq x\underbrace{1\cdots 1}_{c}\underbrace{1\cdots 1}_{a-c-1}\underbrace{0\cdots 0}_{b-c-1}\text{ and }y^{\prime}\coloneqq y\underbrace{1\cdots 1}_{c}\underbrace{0\cdots 0}_{a-c-1}\underbrace{1\cdots 1}_{b-c-1}\enspace.$
Thus, $\operatorname{\mathsf{DISJ}}_{n^{\prime}}^{1}$ is a submatrix of $f$. By 4.4, we have
@@
    $x^{\prime}\coloneqq x\overline{x}\underbrace{0\cdots 0}_{k}\underbrace{1\cdots 1}_{k^{\prime}}\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{n-b-2k}\text{ and }y^{\prime}\coloneqq y\underbrace{0\cdots 0}_{k}\overline{y}\underbrace{1\cdots 1}_{k^{\prime}}\underbrace{1\cdots 1}_{b-a}\underbrace{0\cdots 0}_{n-b-2k}\enspace.$
    Observe that $x^{\prime},y^{\prime}\in\{0,1\}^{n}$ and $|x^{\prime}|=a,|y^{\prime}|=b$. Moreover, $g(x,y)=D(|x\land y|)$ for predicate $D:\left\{0,1,...,k\right\}\rightarrow\left\{-1,1\right\}$ such that $D(z)=f_{a,b}(z+k^{\prime})$ for every $z\in[0,k]$. Thus, we have $D(c-k^{\prime})\neq D(c-k^{\prime}+1)$. By 4.3, we have
    $\log\operatorname{\mathrm{rank}}(g)=\Omega\left(\log\left(\sum_{i=c-k^{\prime}+1}^{k}\binom{k}{i}\right)\right)\enspace.$
@@
    $\operatorname{\mathsf{DISJ}}_{m}^{l-c}(x,y)=f_{a,b}(|x^{\prime}\land y^{\prime}|)=f(x^{\prime},y^{\prime})\enspace,$
@@
    $x^{\prime}\coloneqq x\underbrace{1\cdots 1}_{c}\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{l^{\prime}}\underbrace{1\cdots 1}_{l^{\prime}}\text{ and }y^{\prime}\coloneqq y\underbrace{1\cdots 1}_{c}\underbrace{1\cdots 1}_{b-a}\underbrace{1\cdots 1}_{l^{\prime}}\underbrace{0\cdots 0}_{l^{\prime}}.$
    Thus, $\operatorname{\mathsf{DISJ}}_{m}^{l-c}$ is a submatrix of $f$. By 4.4, we have
    $\log\operatorname{\mathrm{rank}}\left(f\right)\geq\log\operatorname{\mathrm{rank}}(\operatorname{\mathsf{DISJ}}_{m}^{l-c})=\Omega\left(\log\binom{m}{l-c}\right)=\Omega\left(l-c\right)=\Omega\left(a\right)\enspace.$
@@
    $x^{\prime}\coloneqq x\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{a-c}\text{ and }y^{\prime}\coloneqq y\underbrace{1\cdots 1}_{b-a}\underbrace{0\cdots 0}_{a-c}\enspace.$
    Thus, $\operatorname{\mathsf{EQ}}_{n-b+c}^{c}$ is a submatrix of $f$. By 4.4, we have
    $\log\operatorname{\mathrm{rank}}\left(f\right)\geq\log\operatorname{\mathrm{rank}}(\operatorname{\mathsf{EQ}}_{n-b+c}^{c})=\Omega\left(\log\binom{n-b+c}{c}\right)=\Omega\left(c\right)=\Omega\left(a\right)\enspace.$
@@ 5 Log-Approximate-Rank Conjecture for Permutation-Invariant Functions
We discuss Theorem 1.7. In particular, we use the following two lemmas (proved in Appendix D) to prove Theorem 1.7. Additionally, we note that Lemmas 5.1 and 5.2 are variants of Lemmas 3.7 and 3.8.
@@
    $\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\right)$ for $\ell_{1},\ell_{2},\ell_{3}\geq 0$ such that $\ell_{1}+\ell_{2}+\ell_{3}\leq\ell$;
@@
    $\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)=\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$; and
@@
    $\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}\right)$ for $k\geq 1$.
@@
Fix $k\in\mathbb{Z}$. Let $l$ be a half-integer such that $l\leq k/2$. We have
$\log\left(\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)=\Omega\left(\sqrt{kl}\right).$
@@
$\log\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)\enspace.$
Since $\mathsf{ESetInc}_{a,b,c,g}^{n}$ is a subfunction of $f$, we have

### v2

@@ 3.2 Randomized and Quantum Communication Complexity Upper Bound
      $p=\frac{|x\land y|}{n-(a+b)+2|x\land y|}=\frac{1}{\frac{n-(a+b)}{|x\land y|}+2}.$
      Notice that $p$ is an increasing function of $|x\land y|$. As a result, if $|x\land y|\leq c-g$, then $p\leq\frac{c-g}{m-2g}$; if $|x\land y|\geq c+g$, then $p\geq\frac{c+g}{m+2g}$. Let
      $\beta$ $\coloneqq\frac{1}{2}\left(\frac{c+g}{m+2g}+\frac{c-g}{m-2g}\right)$
      $=\frac{cm-2g^{2}}{m^{2}-4g^{2}}$
      $=O\left(\frac{c}{m}\right),$
      $\epsilon$ $\coloneqq\frac{1}{2}\left(\frac{c+g}{m+2g}-\frac{c-g}{m-2g}\right)$
      $=\frac{gm}{m^{2}-4g^{2}}$
      $=\Omega\left(\frac{g}{m}\right).$
      For any $x\in\{0,1\}^{n}$, we let $S_{x}\coloneqq\left\{i:x_{i}=1\right\}$. By 3.11, Alice and Bob can sample $i$ from $S_{\overline{x}\oplus y}$ uniformly using $O(\log n)$ bits communication. Since $i\in S_{\overline{x}\oplus y}$, if $x_{i}=y_{i}=1$, then $i\in S_{x\land y}$; if $x_{i}=y_{i}=0$, then $i\notin S_{x\land y}$. By 3.10, using $O\left(\frac{\beta}{\epsilon^{2}}\right)=O\left(\frac{mc}{g^{2}}\right)=O\left(\frac{n_{1}n_{2}}{g^{2}}\right)$ samples, Alice and Bob can decide whether $p\leq\beta-\epsilon$ or $p\geq\beta+\epsilon$ with bounded error. Equivalently, Alice and Bob can distinguish $|x\land y|\leq c-g$ from $|x\land y|\geq c+g$ with bounded error. By error reduction, using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log\log n\right)$ samples, they can decide whether $|x\land y|\geq c-g$ or $|x\land y|\leq c+g$ with success probability at least $1-1/(6\log n)$. Thus, the communication complexity is $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$.
@@
      Case 3.2: $a+b=n$. Alice and Bob generate new inputs $x^{\prime}=x0$ and $y^{\prime}=y0$ (pad a zero after the original input). Then we have
      $\mathrm{SetInc}_{a,b,c,g}^{n}(x,y)=\mathrm{SetInc}_{a,b,c,g}^{n+1}(x^{\prime},y^{\prime}).$
@@ 4 Log-Rank Conjecture for Permutation-Invariant Functions
Theorem 1.9 states the Log-rank Conjecture for permutation-invariant functions. We argue for the lower bound (Lemma 4.1) and the upper bound (Lemma 4.4) separately.
Lemma 4.1 (Lower Bound).
Fix $n\in{\mathbb{Z}^{+}}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be a non-trivial total permutation-invariant function. For every $a,b\in[n]$ such that $f_{a,b}$ is not a constant function, we have
@@
where $f_{a,b}$ satisfies $f_{a,b}(|x\land y|)=f(x,y)$ for $x,y\in\{0,1\}^{n}$.
Proof.
@@
Fact 4.2 ([15], merging Corollary 6 with Lemma 4).
Fix $n\in{\mathbb{Z}^{+}}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be defined as $f(x,y)\coloneqq D(|x\land y|)$ for some predicate $D:\left\{0,1,...,n\right\}\rightarrow\left\{-1,1\right\}$. If $t$ is the smallest integer such that $D(t)\neq D(t-1)$, then $\log\operatorname{\mathrm{rank}}(f)=\Omega\left(\log\left(\sum_{i=t}^{n}\binom{n}{i}\right)\right)$.
Fact 4.3.
Fix $n\in\mathbb{Z}^{+}$. Let $\mathcal{X},\mathcal{Y}\coloneqq\left\{x\in\{0,1\}^{n}:|x|=k\right\}$, where $k\leq n/2$. Let $\operatorname{\mathsf{DISJ}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{-1,1\right\}$ and $\operatorname{\mathsf{EQ}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{-1,1\right\}$ be defined as
$\operatorname{\mathsf{DISJ}}_{n}^{k}(x,y)$ $1 if |xy|0,\displaystyle\coloneqq$
@@
We refer interesting readers to Section A.2 for the proof of 4.3. Now we prove the lemma by casing on the values of $a$ and $b$.
@@
$\operatorname{\mathsf{DISJ}}_{n^{\prime}}^{1}(x,y)=f_{a,b}(|x\land y|+c)=f(x^{\prime},y^{\prime}),$
@@
[…]
Thus, $\operatorname{\mathsf{DISJ}}_{n^{\prime}}^{1}$ is a submatrix of $f$. By 4.3, we have
@@
    $x^{\prime}$ $\coloneqq x\overline{x}\underbrace{0\cdots 0}_{k}\underbrace{1\cdots 1}_{k^{\prime}}\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{n-b-2k}$
    $y^{\prime}$ $\coloneqq y\underbrace{0\cdots 0}_{k}\overline{y}\underbrace{1\cdots 1}_{k^{\prime}}\underbrace{1\cdots 1}_{b-a}\underbrace{0\cdots 0}_{n-b-2k}.$
    Observe that $x^{\prime},y^{\prime}\in\{0,1\}^{n}$ and $|x^{\prime}|=a,|y^{\prime}|=b$. Moreover, $g(x,y)=D(|x\land y|)$ for predicate $D:\left\{0,1,...,k\right\}\rightarrow\left\{-1,1\right\}$ such that $D(z)=f_{a,b}(z+k^{\prime})$ for every $z\in[0,k]$. Thus, we have $D(c-k^{\prime})\neq D(c-k^{\prime}+1)$. By 4.2, we have
    $\log\operatorname{\mathrm{rank}}(g)=\Omega\left(\log\left(\sum_{i=c-k^{\prime}+1}^{k}\binom{k}{i}\right)\right)$.
@@
    $\operatorname{\mathsf{DISJ}}_{m}^{l-c}(x,y)=f_{a,b}(|x^{\prime}\land y^{\prime}|)=f(x^{\prime},y^{\prime}),$
@@
    $x^{\prime}$ $\coloneqq x\underbrace{1\cdots 1}_{c}\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{l^{\prime}}\underbrace{1\cdots 1}_{l^{\prime}},$
    $y^{\prime}$ $\coloneqq y\underbrace{1\cdots 1}_{c}\underbrace{1\cdots 1}_{b-a}\underbrace{1\cdots 1}_{l^{\prime}}\underbrace{0\cdots 0}_{l^{\prime}}.$
    Thus, $\operatorname{\mathsf{DISJ}}_{m}^{l-c}$ is a submatrix of $f$. By 4.3, we have
    $\log\operatorname{\mathrm{rank}}\left(f\right)$ $\geq\log\operatorname{\mathrm{rank}}(\operatorname{\mathsf{DISJ}}_{m}^{l-c})$
    $=\Omega\left(\log\binom{m}{l-c}\right)$
    $=\Omega\left(l-c\right)$
    $=\Omega\left(a\right).$
@@
    $x^{\prime}\coloneqq x\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{a-c}\text{ and }y^{\prime}\coloneqq y\underbrace{1\cdots 1}_{b-a}\underbrace{0\cdots 0}_{a-c}.$
    Thus, $\operatorname{\mathsf{EQ}}_{n-b+c}^{c}$ is a submatrix of $f$. By 4.3, we have
    $\log\operatorname{\mathrm{rank}}\left(f\right)$ $\geq\log\operatorname{\mathrm{rank}}(\operatorname{\mathsf{EQ}}_{n-b+c}^{c})$
    $=\Omega\left(\log\binom{n-b+c}{c}\right)$
    $=\Omega\left(c\right)$
    $=\Omega\left(a\right).$
Lemma 4.4 (Upper Bound).
Fix $n\in{\mathbb{Z}^{+}}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be a non-trivial total permutation-invariant function. Then $D(f)$ is
$O\left(\max_{a,b\in[n]:f_{a,b}\text{ is not constant}}\min\left\{a,b,n-a,n-b\right\}\cdot\log n\right),$
where $f_{a,b}$ satisfies $f_{a,b}(|x\land y|)=f(x,y))$ for $x,y\in\{0,1\}^{n}$.
Proof.
$\log\binom{n}{a}\leq\min\left\{a,n-a\right\}\cdot\log n.$ (7)
Combining (5), (6), and (7), the communication cost of the protocol is
@@ 5 Log-Approximate-Rank Conjecture for Permutation-Invariant Functions
We discuss Theorem 1.10. In particular, we use the following two lemmas (proved in Appendix C) to prove Theorem 1.10. Additionally, we note that Lemmas 5.1 and 5.2 are variants of Lemmas 3.6 and 3.7.
@@
    $\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)\leq\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\right)$ for $\ell_{1},\ell_{2},\ell_{3}\geq 0$ such that $\ell_{1}+\ell_{2}+\ell_{3}\leq\ell$;
@@
    $\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)=\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{a,n-b,a-c,g}^{n}\right)=\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{n-a,b,b-c,g}^{n}\right)$; and
@@
    $\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)\leq\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{ka,kb,kc,kg}^{kn}\right)$ for $k\geq 1$.
@@
Fix $k\in\mathbb{Z}$. Let $l$ be a half-integer such that $0<l\leq k/2$. We have
$\log\left(\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)=\Omega\left(\sqrt{kl}\right).$
@@
$\log\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
Since $\mathrm{ESetInc}_{a,b,c,g}^{n}$ is a subfunction of $f$, we have